OpenStudy (anonymous):

Which is the simplified form of 18 times q to the fourth power times r all over 6 times q cubed times r to the fifth power.? t0mmiere can you help 3 times r to the fourth power all over q. 12 times r to the fourth power all over. 3 times q all over r to the fourth power. 12 times q all over r to the fourth power.

OpenStudy (zzr0ck3r):

\[\frac{18q^4r}{6q^3r^5} =\frac{18}{6}q^{4-3}r^{1-4}=3qr^{-4}=\frac{3q}{r^4}\]

OpenStudy (anonymous):

thank you zzr0ck :)

OpenStudy (anonymous):

can you help me with one more Simplify the expression. x times the square root of the quantity 64 x cubed, plus 2 x times the square root of x, plus 8 times the square root of the quantity 2 x

OpenStudy (zzr0ck3r):

\[x*\sqrt{64x^3}+2x*\sqrt{x}+8*\sqrt{2x}?\]

OpenStudy (anonymous):

yes thats it

OpenStudy (zzr0ck3r):

\[8*x^2\sqrt{x}+ 2x \sqrt{x}+ 8*\sqrt{2}*\sqrt{x}=\sqrt{x}*(8x^2+2x+8\sqrt{2})\]

OpenStudy (anonymous):

it also askes what the leading coefficient of it would be when it is simplified would it be the eight or sqrt of x

OpenStudy (anonymous):

i think its 8x^2 am i right?

OpenStudy (zzr0ck3r):

8

OpenStudy (anonymous):

ok thank you again:)

OpenStudy (zzr0ck3r):

np